Solve using the principle of zero products.
(x - 2)(2x + 8) = 0

Answer by Earlsdon(6294)

The "zero products" principle means: If a*b = 0, then either a = 0, b = 0, or both, so, if...
(x-2)(2x+8) = 0 then;
x-2 = 0 or 2x+8 = 0 and...
x = 2 or 2x = -8 so x = -4
The solutions are:
x = 2
x = -4